OpenStudy (anonymous):

Brian wants to have $7000 in the bank in 10 years. He deposits $3000 today at 8% interest compounded semiannually. How much additional money will he need to reach the desired $7000

OpenStudy (campbell_st):

well you need to find the principal using the compound interest formula \[A = P \times(1 + \frac{r}{100})^n\] so in your question, you know A = 7000 and r = 8 and the number of time periods n = 20 (semi annually means twice per year) so your formula is \[7000 = P \times (1 + \frac{8}{100})^{20}\] now you need to solve for P

OpenStudy (campbell_st):

hope it helps

OpenStudy (amistre64):

is he going to make payments? or just needing to get more money to deposit all at once?